# TopoTune: a framework for generalized combinatorial complex neural networks

### v2

@@ 6 Experiments
We present experiments showcasing a broad class of GCCN’s constructed with TopoTune. These models consistently match, outperform, or improve upon existing CCNNs, often with smaller model sizes. TopoTune’s integration into the TopoBenchmark experiment infrastructure ensures a fair comparison with CCNNs from the literature, as implementations of data processing, domain lifting, and training are homogeonized. Moreover, TopoBenchmark aggregates TDL’s leading open-source models from TopoX Hajij et al. (2024b), making it an ideal framework for objectively evaluating GCCNs against the current state of the field.
6.1 Experimental Setup
We generate our library by considering ten possible choices of neighborhood structure $\mathcal{N}_{\mathcal{C}}$ (including both regular and per-rank, see Appendix C.1) and five possible choices of $\omega_{\mathcal{N}}$: GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2017), GIN (Xu et al., 2019b), GraphSAGE (Hamilton et al., 2017), and Transformer (Vaswani et al., 2017). We import these models directly from PyTorch Geometric (Fey & Lenssen, 2019) and PyTorch Paszke et al. (2019). TopoTune enables running GCCNs on both graph expansions previously discussed: an ensemble of strictly augmented Hasse graphs (eq. 5) and a single augmented Hasse graph (eq. 4). In the latter case, each GCCN layer simply contains a single $\omega_{\mathcal{N}}$ (e.g. GNN or Transformer) which processes the single augmented Hasse graph.
Datasets.
We include a wide range of tasks (see Appendix C.2). MUTAG, PROTEINS, NCI01, and NCI09 (Morris et al., 2020) are graph-level classification tasks which seek to predict some property about molecules or proteins. ZINC (Irwin et al., 2012) aims to predict a graph-level property related to molecular solubility through regression. At the node level, the Cora, CiteSeer, and PubMed tasks (Yang et al., 2016) involve classifying publications (nodes) within citation networks. We consider two particular cases of combinatorial complexes, simplicial and cellular complexes, leveraging TopoBenchmark’s data lifting processes to infer higher-order relationships in these datasets (see Appendix C.2 for further details).
6.2 Results and Discussion
@@
Table 1 portrays a cross-comparison between top-performing CCNN models and our class of GCCNs. GCCNs outperform CCNNs in the simplicial and cellular domains across all datasets. Notably, GCCNs in these domains achieve comparable results to hypergraph CCNNs, a feat unattainable by existing CCNNs in node-level tasks. Moreover, we find both of GCCN’s underlying notions (see Section 4) to be advantageous: (i) Table 1 shows that representing a complex as an ensemble of augmented Hasse graphs consistently yields better performance than using a single augmented Hasse graph. (ii) Some GCCNs with a per-rank neighborhood structure outperform not only CCNNs but also all other GCCNs using regular neighborhoods. For example, this is the case for a cellular GCCN with $\omega_{\mathcal{N}}$ on MUTAG. Its lightweight, per-rank neighborhood structure makes it 19% the size of the best cellular CCNN on this task.
GCCNs are smaller than CCNNs.
GCCNs are generally more parameter efficient than existing CCNNs in simplicial and cellular domains, and in some instances (MUTAG, NCI1, NCI09), even surpass hypergraph CCNNs in size efficiency. Even as GCCNs become more resource-intensive for large graphs with high-dimensional embeddings—as seen in node-level tasks—they maintain a competitive edge. For instance, on the Citeseer dataset, a GCCN ($\omega_{\mathcal{N}}$ = GraphSAGE) outperforms the best existing CCNN while being 28% smaller. We refer to Table 4 in Appendix D.
GCCNs improve existing CCNNs.
@@
TopoTune makes it easy to iterate upon and improve preexisting CCNNs by replicating their architecture in a GCCN setting. For example, TopoTune can generate a counterpart GCCN by replicating a CCNN’s neighborhood structure, aggregation, and training scheme. We show in Table 2 that counterpart GCCNs often achieve comparable or better results than SCCN (Yang et al., 2022) and CWN (Bodnar et al., 2021a) just by sweeping over additional choices of $\omega_{\mathcal{N}}$ (same as in Table 1) available in PyTorch Geometric (Fey & Lenssen, 2019). In the single augmented Hasse graph regime, GCCN models are consistently more lightweight, up to half their size (see Table 5 in Appendix D for details).
Model MUTAG PROTEINS NCI1 NCI109 Cora Citeseer PubMed
SCCN Yang et al. (2022)
Benchmark results Telyatnikov et al. (2024) 70.64 ± 5.90 74.19 ± 2.86 76.60 ± 1.75 77.12 ± 1.07 82.19 ± 1.07 69.60 ± 1.83 88.18 ± 0.32
CWN Bodnar et al. (2021a)
Benchmark results Telyatnikov et al. (2024) 80.43 ± 1.78 76.13 ± 2.70 73.93 ± 1.87 73.80 ± 2.06 86.32 ± 1.38 75.20 ± 1.82 88.64 ± 0.36
Performance-Cost Tradeoff.
@@
By easily sweeping over a wide landscape of possible GCCNs for a given task, TopoTune can help identify “lottery-ticket" models that maximize performance while minimizing computational cost. Fig. 5 illustrates this concept by comparing the performance and size of selected GCCNs across three datasets (see Appendix E for plots on all datasets). On the PROTEINS dataset, two GCCNs using per-rank neighborhood structures (orange and purple) achieve performance within 2% of the best result, while using as little as 48% of the parameters. Similarly, on ZINC, lightweight neighborhood structures (orange and dark green) deliver strong results with reduced parameter costs. We remark that node-level tasks do not always benefit from this parameter reduction. For instance, Citeseer maintains a large model size regardless of neighborhood structure, primarily due to its large graph size and high-dimensional input features. Fig. 5 also provides insights into the impact of neighborhood message functions. On ZINC, GIN clearly outperforms all other models. In less clear-cut cases, we observe a trade-off between neighborhood structure and message function complexity. While GIN and GraphSAGE are generally more parameter-intensive than GAT and GCN, we find that more complex base models on lightweight neighborhood structures perform comparably to simpler base models on more complete neighborhood structures.
Figure 5: GCCN task-specific performance versus size. We compare relative performance and parameter size of various GCCNs across three datasets on the cellular domain, two graph-level (left, middle) and one node-level (right). Each point represents a different combination of neighborhood structure $\mathcal{N}_{\mathcal{C}}$ and message function $\omega_{\mathcal{N}}$. The amount of layers is kept constant according to the best performing model. Both the x-axis and y-axis are scaled relative to the best performing model.
@@ 7 Conclusion
This work introduces a simple yet powerful graph-based methodology for constructing Generalized Combinatorial Complex Neural Networks (GCCNs), TDL architectures that generalize and subsume standard CCNNs. Additionally, we introduce TopoTune, the first lightweight software module for systematically and easily implementing new TDL architectures, all the while leveraging existing GNN (or any other models’) software. Building upon this foundation, we envision several promising directions for future research. Application-specific TDL could seek to customize GCCNs for potentially sparse or multimodal datasets. Additionally, TopoTune’s graph-based approach paves the way for extending state-of-the-art GNN architectures beyond those currently implemented in open-source libraries such as PyTorch Geometric. We hope TopoTune will not only accelerate TDL research but also help bridge the gap with other machine learning fields.
Appendix A Proofs
A.1 Proof of Generality
The proof is trivial. It is sufficient to set $\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{N}},\mathcal{G}_{\mathcal{N}})$ to $\{\bigoplus_{y\in\mathcal{N}(\sigma)}\psi_{\mathcal{N},\textrm{rk}(\sigma)}\left(\mathbf{h}^{l}_{\sigma},\mathbf{h}^{l}_{\tau}\right)\}_{\sigma\in\mathcal{C}}$ in (8) as all $y\in\mathcal{N}(\sigma)$ are part of the node set $\mathcal{C}_{\mathcal{N}}$ of the strictly augmented Hasse graph of $\mathcal{N}$ by definition.
A.2 Proof of Equivariance
@@
Proof of Proposition 2. We follow the approach from (Bodnar et al., 2021a). Given any permutation matrix $\mathbf{P}$, for a cell $\sigma_{i}$, let us denote its permutation as $\sigma_{\mathbf{P}(i)}$ with an abuse of notation. Let $\mathbf{h}^{l+1}_{\sigma_{i}}$ be the output embedding of cell $\sigma_{i}$ for the $l$-th layer of a GCCN taking $(\mathbf{H}^{l},\{\mathbf{N}_{\mathcal{N}}\}_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}})$ as input, and $\mathbf{h}^{l+1}_{\sigma_{\mathbf{P}(i)}}$ be the output embedding of cell $\sigma_{\mathbf{P}(i)}$ for the same GCCN layer taking $\left(\mathbf{P}\mathbf{H}^{l},\{\mathbf{PN}_{\mathcal{N}}\mathbf{P}^{T}\}_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}\right)$ as input. To prove the permutation equivariance, it is sufficient to show that $\mathbf{h}^{l+1}_{\sigma_{i}}=\mathbf{h}^{l+1}_{\sigma_{\mathbf{P}(i)}}$ a the update function $\phi$ is row-wise, i.e. it independetly act on each cell. To do so, we show that the (multi-)set of embeddings being passed to the neighborhood message function, aggregation, and update functions are the same for the two cells $\sigma_{i}$ and $\sigma_{\mathbf{P}(i)}$. The neighborhood message functions act on the strictly augmented Hasse graph of $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$ of $\mathcal{N}$, thus we work with the submatrix $\widetilde{\mathbf{N}}_{\mathcal{N}}$. The neighborhood message function is assumed to be node permutation equivariant, i.e., denoting again the embeddings of the cells in $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$ with $\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}}\in\mathbb{R}^{|\mathcal{C}_{\mathcal{N}}|\times F^{l}}$ and identifying $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$ with $\widetilde{\mathbf{N}}_{\mathcal{N}}$, it holds that $\omega_{\mathcal{N}}(\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\widetilde{\mathbf{N}}_{\mathcal{N}}\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}^{T})=\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\widetilde{\mathbf{N}}_{\mathcal{N}})$, where $\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}$ is the submatrix of $\mathbf{P}$ given by the rows and the columns corresponding to the cells in $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$. This assumption, together with the assumption that the inter-neighborhood aggregation is assumed to be cell permutation invariant, i.e. $\bigotimes_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\widetilde{\mathbf{N}}_{\mathcal{N}})=\bigotimes_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\widetilde{\mathbf{N}}_{\mathcal{N}})$, trivially makes the overall composition of the neighborhood message function with the inter-neighborhood aggregation cell permutation invariant. This fact, together with the fact that the (labels of) the neighbors of the cell $\sigma_{i}$ in $\mathcal{N}$ are given by the nonzero elements of the $i$-th row of $\mathbf{N}_{\mathcal{N}}$, or the corresponding row of $\widetilde{\mathbf{N}}_{\mathcal{N}}$, and that the columns and rows of $\widetilde{\mathbf{N}}_{\mathcal{N}}$ are permuted in the same way the rows of the feature matrix $\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}}$ are permuted, implies

### v3

@@ 6 Experiments
We present experiments showcasing a broad class of GCCN’s constructed with TopoTune. These models consistently match, outperform, or finetune existing CCNNs, often with smaller model sizes. TopoTune’s integration into the TopoBenchmark experiment infrastructure ensures a fair comparison with CCNNs from the literature, as data processing, domain lifting, and training are homogeonized.
6.1 Experimental Setup
We generate our class of GCCNs by considering ten possible choices of neighborhood structure $\mathcal{N}_{\mathcal{C}}$ (including both regular and per-rank, see Appendix E.1) and five possible choices of $\omega_{\mathcal{N}}$: GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2017), GIN (Xu et al., 2019a), GraphSAGE (Hamilton et al., 2017), and Transformer (Vaswani et al., 2017). We import these models directly from PyTorch Geometric (Fey & Lenssen, 2019) and PyTorch (Paszke et al., 2019). TopoTune enables running GCCNs on both an ensemble of strictly augmented Hasse graphs (eq. 6) and an augmented Hasse graph (eq. 5). While CCNN results reflect extensive hyperparameter tuning by Telyatnikov et al. (2024), we largely fix GCCN training hyperparameters using the TopoBenchmark default configuration.
Datasets.
We include a wide range of benchmark tasks (see Appendix E.2). MUTAG, PROTEINS, NCI01, and NCI09 (Morris et al., 2020) are graph-level classification tasks about molecules or proteins. ZINC (Irwin et al., 2012) (subset) is a graph-level regression task about solubility. At the node level, the Cora, CiteSeer, and PubMed tasks (Yang et al., 2016) involve classifying publications within citation networks. We consider two topological domains: simplicial and cellular complexes. We use TopoBenchmark’s lifting processes to infer higher-order relationships in these data. We use node features to construct edge and face features.
6.2 Results and Discussion
Figure 5: GCCN performance versus size. We compare various GCCNs across three datasets on the cellular domain, two graph-level (left, middle) and one node-level (right). Each GCCN (point) has a different neighborhood structure $\mathcal{N}_{\mathcal{C}}$, some of which can only be represented as per-rank structures ($\square$ in legend), and message function $\omega_{\mathcal{N}}$. The amount of layers is kept constant according to the best performing model. The axes are scaled relative to this model.
@@ 6.2 Results and Discussion
Table 1 portrays a cross-comparison between top-performing CCNN models and our class of GCCNs. GCCNs outperform CCNNs in the simplicial and cellular domains in all datasets. Notably, GCCNs in these domains achieve comparable results to hypergraph CCNNs, a feat unattainable by CCNNs in node-level tasks. Out of the 16 domain/dataset combinations considered in our experiments, GCCNs outperform the best counterpart CCNN by $>1\sigma$ in 11 cases. Representing complexes as ensembles of augmented Hasse graphs, rather than a single augmented Hasse graph, consistently improves results.
Generalizing existing CCNNs to GCCNs improves performance.
@@
TopoTune makes it easy to iterate upon and improve preexisting CCNNs by replicating their architecture in a GCCN setting. For example, TopoTune can generate a counterpart GCCN by replicating a CCNN’s neighborhood structure, aggregation, and training scheme. We show in Table 3 in Appendix F that counterpart GCCNs can achieve comparable or better results than SCCN (Yang et al., 2022) and CWN (Bodnar et al., 2021a) just by sweeping over additional choices of $\omega_{\mathcal{N}}$. In the single augmented Hasse graph regime, GCCN models are consistently more lightweight, up to half their size (see Table 5).
GCCNs perform competitively to CCNNs with fewer parameters.
GCCNs are often more parameter efficient than existing CCNNs in simplicial and cellular domains, and in some instances (MUTAG, NCI1, NCI09), even in the hypergraph domain. We refer to Table 4. Even as GCCNs become more parameter-intensive for large graphs with high-dimensional embeddings—as seen in node-level tasks—they remain competitive. For instance, on the Citeseer dataset, a GCCN ($\omega_{\mathcal{N}}$ = GraphSAGE) outperforms the best existing CCNN while being 28% smaller. Training times in Appendix H show GCCNs train comparably on smaller datasets but slow down on larger ones, likely due to TopoTune’s on-the-fly graph expansion. Preprocessing this expansion in future work could mitigate the lag.
TopoTune finds parameter-efficient GCCNs.
@@
By easily exploring a wide landscape of possible GCCNs for a given task, TopoTune helps identify models that maximize performance while minimizing model size. Fig. 5 illustrates this trade-off by comparing the performance and size of selected GCCNs (see Appendix I for more). On the PROTEINS dataset, two GCCNs using per-rank neighborhood structures (orange and purple) achieve performance within 2% of the best result while requiring as little as 48% of the parameters. This reduction is due to fewer neighborhoods $\mathcal{N}$, resulting in fewer $\omega_{\mathcal{N}}$ blocks per GCCN layer. Similarly, on ZINC, lightweight neighborhood structures (orange and dark green) are competitive with reduced parameter costs. Node-level tasks, see less benefit, likely due to the larger graph sizes and higher-dimensional input features.
Impactfulness of GNN choice is dataset specific.
Fig. 5 also provides insights into the impact of neighborhood message functions. On ZINC, GIN clearly outperforms all other models, which do not even appear in the plot’s range. In the less clear-cut cases of PROTEINS and Citeseer, we observe a trade-off between neighborhood structure and message function complexity. We find that larger base models (GIN, GraphSAGE) on lightweight neighborhood structures perform comparably to simpler base models (GAT, GCN) on larger neighborhood structures. This tradeoff warrants further research on the dataset-specific importance of neighborhood choice, or lack thereof.
@@ 7 Conclusion
This work introduces a simple yet powerful graph-based methodology for constructing Generalized Combinatorial Complex Neural Networks (GCCNs), TDL architectures that generalize and subsume standard CCNNs. Additionally, we introduce TopoTune, the first lightweight software module for systematically and easily implementing new TDL architectures across topological domains. In doing so, we have addressed, either in part or in full, 7 of the 11 open problems of the field defined by some of its leaders in Papamarkou et al. (2024). Future work includes customizing GCCNs for application-specific and potentially sparse or multimodal datasets, and leveraging software from state-of-the-art GNNs. TopoTune will also help bridge the gap with other fields such as attentional learning and $k$-hop higher-order GNNs (Morris et al., 2019; Maron et al., 2019).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Domains of Topological Deep Learning
Figure 6: Topological Deep Learning Domains. Nodes in blue, (hyper)edges in pink, and faces in dark red. Figure adopted from Papillon et al. (2023).
We summarize the different discrete domains leveraged within TDL and, in doing so, contextualize how combinatorial complexes generalize all of them. To that end, we will closely follow the description of (Papillon et al., 2023), using as well its very clarifying Figure 6. We recommend this survey for a high-level overview of TDL literature, and the more extensive work of (Hajij et al., 2023) for a detailed formulation of the field. We also refer to Appendix C of Battiloro et al. (2025) for a concise mathematical description of each domain. From left to right in Figure 6, the different domains in TDL are:
Traditional Discrete Domains
Set / Pointcloud.
A collection of points called nodes without any additional structure.
Graph.
A set of points (nodes) connected with edges that denote pairwise relationships.
Set + Part-Whole Relations
Simplicial Complex.
A generalization of a graph that incorporates hierarchical part-whole relations through the multi-scale construction of cells. Nodes are rank 0-cells that can be combined to form edges (rank 1 cells). Edges are, in turn, combined to form faces (rank 2 cells), which are combined to form volumes (rank 3 cells), and so on. In particular, each cell $\sigma$ in a simplicial complex must contain all lower dimensional cells $\tau$ such that $\tau\subseteq\sigma$. Therefore, faces must be triangles, volumes must be tetrahedrons, and so forth.
Cellular Complex.
A generalization of an simplicial complex in which cells are not limited to simplexes, but may instead take any shape: faces can involve more than three nodes, volumes more than four faces, and so on. This flexibility endows CCs with greater expressivity than simplicial complexes (Bodnar et al., 2021a), but still edges only connect pairs of nodes.
Set + Set-Type Relations
Hypergraph:
A generalization of a graph, in which higher-order edges called hyperedges can connect arbitrary sets of two or more nodes. Connections in HGs represent set-type relationships, in which participation in an interaction is not implied by any other relation in the system. This makes HGs an ideal choice for data with abstract and arbitrarily large interactions of equal importance, such as semantic text and citation networks.
Set + Part-Whole and Set-Type Relations
Combinatorial Complex:
A structure that combines features of hypergraphs and cellular complexes. Like a hypergraph, edges may connect any number of nodes. Like a cellular complex, cells can be combined to form higher-ranked structures. Hence, combinatorial complexes generalize all other topological domains.
Appendix B Proofs
B.1 Proof of Generality
The proof is straightforward. It is sufficient to set $\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{N}},\mathcal{G}_{\mathcal{N}})$ to $\{\bigoplus_{y\in\mathcal{N}(\sigma)}\psi_{\mathcal{N},\textrm{rk}(\sigma)}\left(\mathbf{h}^{l}_{\sigma},\mathbf{h}^{l}_{\tau}\right)\}_{\sigma\in\mathcal{C}}$ in (8) as all $y\in\mathcal{N}(\sigma)$ are part of the node set $\mathcal{C}_{\mathcal{N}}$ of the strictly augmented Hasse graph of $\mathcal{N}$ by definition.
B.2 Proof of Equivariance
@@
Proof of Proposition 4.2. We follow the approach from (Bodnar et al., 2021a). Given any permutation matrix $\mathbf{P}$, for a cell $\sigma_{i}$, let us denote its permutation as $\sigma_{\mathbf{P}(i)}$ with an abuse of notation. Let $\mathbf{h}^{l+1}_{\sigma_{i}}$ be the output embedding of cell $\sigma_{i}$ for the $l$-th layer of a GCCN taking $(\mathbf{H}^{l},\{\mathbf{N}_{\mathcal{N}}\}_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}})$ as input, and $\mathbf{h}^{l+1}_{\sigma_{\mathbf{P}(i)}}$ be the output embedding of cell $\sigma_{\mathbf{P}(i)}$ for the same GCCN layer taking $\left(\mathbf{P}\mathbf{H}^{l},\{\mathbf{PN}_{\mathcal{N}}\mathbf{P}^{T}\}_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}\right)$ as input. To prove the permutation equivariance, it is sufficient to show that $\mathbf{h}^{l+1}_{\sigma_{i}}=\mathbf{h}^{l+1}_{\sigma_{\mathbf{P}(i)}}$ as the update function $\phi$ is row-wise, i.e., it independently acts on each cell. To do so, we show that the (multi-)set of embeddings being passed to the neighborhood message function, aggregation, and update functions are the same for the two cells $\sigma_{i}$ and $\sigma_{\mathbf{P}(i)}$. The neighborhood message functions act on the strictly augmented Hasse graph of $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$ of $\mathcal{N}$, thus we work with the submatrix $\widetilde{\mathbf{N}}_{\mathcal{N}}$. The neighborhood message function is assumed to be node permutation equivariant, i.e., denoting again the embeddings of the cells in $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$ with $\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}}\in\mathbb{R}^{|\mathcal{C}_{\mathcal{N}}|\times F^{l}}$ and identifying $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$ with $\widetilde{\mathbf{N}}_{\mathcal{N}}$, it holds that $\omega_{\mathcal{N}}(\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\widetilde{\mathbf{N}}_{\mathcal{N}}\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}^{T})=\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\widetilde{\mathbf{N}}_{\mathcal{N}})$, where $\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}$ is the submatrix of $\mathbf{P}$ given by the rows and the columns corresponding to the cells in $\mathcal{G}_{\mathcal{C}_{\mathcal{N}}}$. This assumption, together with the assumption that the inter-neighborhood aggregation is assumed to be cell permutation invariant, i.e. $\bigotimes_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}\mathbf{P}_{\mathcal{C}_{\mathcal{N}}}\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\widetilde{\mathbf{N}}_{\mathcal{N}})=\bigotimes_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}\omega_{\mathcal{N}}(\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}},\widetilde{\mathbf{N}}_{\mathcal{N}})$, trivially makes the overall composition of the neighborhood message function with the inter-neighborhood aggregation cell permutation invariant. This fact, together with the fact that the (labels of) the neighbors of the cell $\sigma_{i}$ in $\mathcal{N}$ are given by the nonzero elements of the $i$-th row of $\mathbf{N}_{\mathcal{N}}$, or the corresponding row of $\widetilde{\mathbf{N}}_{\mathcal{N}}$, and that the columns and rows of $\widetilde{\mathbf{N}}_{\mathcal{N}}$ are permuted in the same way the rows of the feature matrix $\mathbf{H}^{l}_{\mathcal{C}_{\mathcal{N}}}$ are permuted, implies